Word problems are really giving me difficulty. Does anyone have a simple easy rule?
d=rt
Kicking my behind.
A jogger runs at a rate of 10mph. He returns, same route at 6mph. Total trip took 1 hour 36 minutes. How far did he run?
Delena
thanks all!

Found 2 solutions by scott8148, richard1234:
Answer by scott8148(6628)

they give you the total time and the rates

as you say, d = rt ___ but also, t = d / r

(d / 10) + (d / 6) = 1.6
___ the rates are in mi per HOUR, so the time needs to be hours; not hours and minutes
___ 36 min = (36 / 60) hr = (6 / 10) hr = .6 hr

multiplying by 30 (LCD) ___ 3d + 5d = 48 ___ 8d = 48

Answer by richard1234(7193)

Suppose the route is d miles long. Using d = rt, we have t = d/r. We have to add two different times (going there, coming back) using this formula. Also, we want consistent units so use 1.6 hours instead of 1 hr 36 minutes.:

Use LCD of 30:





(miles)